### v3

# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

Abstract
Multi-label classification (MLC) often exhibits performance disparities, especially for infrequent or sensitive label categories. We propose FairPO, a novel framework that integrates preference-based loss formulations with group-robust optimization to improve fairness in multi-label classification (MLC), particularly targeting underperforming and sensitive label groups. FairPO partitions labels into a privileged set, targeted for enhanced performance, and a non-privileged set, where baseline performance is maintained. For privileged labels, a preference-based loss, inspired by Direct Preference Optimization (DPO), encourages model scores for true positives to be significantly higher than for their confusing negative counterparts, and scores for true negatives to be significantly lower than for their confusing positive counterparts—addressing hard examples that challenge standard classifiers. For non-privileged labels, a constrained objective ensures performance does not degrade substantially below a reference model. These group-specific objectives are balanced using a Group Robust Preference Optimization (GRPO) formulation, adaptively mitigating bias. Our experiments also include FairPO variants built on recent reference-free preference optimization techniques, namely Contrastive Preference Optimization (CPO) and Simple Preference Optimization (SimPO), which further highlight FairPO’s versatility [^1] Our code is available at GitHub: https://anonymous.4open.science/r/FairPO.
1 Introduction
Multi-label classification (MLC)—assigning a subset of labels from a universe $\mathcal{T}$ to an instance xx—is a pervasive task in fields like image annotation (Wang et al., 2016) and document categorization (Zangari et al., 2024; Schietgat et al., 2010). Typically, models are trained by minimizing per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2014; Sorower, 2010; Tarekegn et al., 2024). However, this global optimization often creates fairness issues (Mehrabi et al., 2022a; Chouldechova, 2017), as imbalanced label frequencies and varying importance lead to disparate performance (Mehrabi et al., 2022b). For instance, a model might excel on common labels while consistently failing on rare but critical ones. This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
Figure 1: FairPO’s methodology
A key challenge lies in the model’s ability to discriminate effectively, especially for hard examples. Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions. For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a confusing negative that the model mistakenly ranks above the true label. Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive). Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2011) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization. Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2024), FairPO recasts parts of the MLC task as learning explicit preferences. We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective. If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied. This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$. If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning. For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$. To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024). The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$. This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting fairness across the labels (see Figure 1).
Our primary contributions are a novel framework for fair MLC using preference signals, a conditional objective targeting hard examples, and a robust optimization strategy to manage fairness trade-offs. Our comprehensive experiments show that FairPO, particularly the CPO-variant, achieves significant gains—for instance, up to 3.44% mAP on the least frequent labels of COCO (Lin et al., 2015)—while robustly maintaining performance on non-privileged labels. We also outline the framework’s applicability to attribute generation in Appendix A.
2 Preliminaries: Preference Optimization Methods
Our framework builds upon recent preference optimization techniques. We briefly review the key formulations.
Direct Preference Optimization (DPO): DPO (Rafailov et al., 2024) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
$h_{\pi_{\theta}}(x,y_{w},y_{l})$ $=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},$ (1)
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(h_{\pi_{\theta}}(x,y_{w},y_{l})\right)\right].$ (2)
where $\sigma$ is the sigmoid function and $\beta$ controls the deviation from $\pi_{\text{ref}}$.
Group Robust Preference Optimization (GRPO): GRPO (Ramesh et al., 2024) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$. Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
$\min_{\pi_{\theta}}\max_{\alpha\in\Delta_{K-1}}\sum_{g=1}^{K}\alpha_{g}L_{\text{Pref}}(\pi_{\theta};\pi_{\text{ref}},D_{g}),$ (3)
where $L_{\text{Pref}}$ is a base preference loss (like $L_{\text{DPO}}$), and $\alpha=(\alpha_{1},...,\alpha_{K})$ are adaptive weights in the probability simplex $\Delta_{K-1}$. The optimization dynamically increases weights $\alpha_{g}$ for groups with higher current loss, focusing learning on the worst-performing groups.
Simple Preference Optimization (SimPO): SimPO (Meng et al., 2024) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$. It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$. It also introduces a target margin $\gamma>0$ into the preference model. The resulting SimPO loss is:
$L_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$ (4)
Contrastive Preference Optimization (CPO): CPO (Xu et al., 2024) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective. It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
$L_{\text{prefer}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)\right]$ (5)
$L_{\text{NLL}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w})\sim D}[\log\pi_{\theta}(y_{w}|x)]$ (6)
$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (7)
This formulation avoids the computational cost of the reference model pass.
3 Methodology: Fair Preference Optimization (FairPO)
3.1 Problem Setup and Fairness Goals
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune parameters $\mathbf{w}_{t}$ for a per-label classifier that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$. Our framework leverages a reference model with parameters $\hat{\mathbf{w}}_{t}$, typically obtained from standard supervised fine-tuning (SFT).
Our framework addresses the problem of group fairness, where the goal is to ensure equitable performance across different predefined groups. In our MLC context, we define these groups not by instance attributes (e.g., demographics), but by partitioning the labels themselves into a privileged set $\mathcal{P}$ (e.g., rare, critical, or historically underperforming labels) and a non-privileged set $\bar{\mathcal{P}}$. Our fairness objective is a nuanced form of equitable treatment: we aim to significantly improve performance for the underperforming privileged group while ensuring that performance for the non-privileged group is not substantially harmed. This goal of targeted improvement without undue harm to other groups is enforced by our robust optimization framework.
For any privileged label $l\in\mathcal{P}$, the model must accurately discriminate its true state from confusing alternatives. For any non-privileged label $j\in\bar{\mathcal{P}}$, performance should not degrade significantly from the reference model. To formalize this, we define two types of confusing sets for a privileged label $l\in\mathcal{P}$: the set of confusing negatives, $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=+1$, and the set of confusing positives, $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=0$. The overall confusing set is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
3.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2024) preference loss. If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$). The DPO term is:
$h_{\mathbf{w}}(x_{i},l,k)=\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right).$ (8)
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
If $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$ is sampled, we prefer $l$ over $k$ (i.e., the true negative $l$ is preferred over the confusing positive $k$, meaning $m(x_{i};\mathbf{w}_{l})$ should be lower than $m(x_{i};\mathbf{w}_{k})$ relative to the reference). The DPO term is (note the swapped order for $k$ and $l$ to reflect $k\succ l$ effectively aiming for $m(x_{i};\mathbf{w}_{l})<m(x_{i};\mathbf{w}_{k})$):
$h_{\mathbf{w}}(x_{i},k,l)=\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right).$ (9)
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
No Confusing Examples ($S_{il}=\emptyset$): We revert to a standard base classification loss (e.g., Binary Cross-Entropy, BCE) for label $l$. This BCE fallback ensures that the model continues to receive learning signals even for labels without any identified confusing counterparts, thereby avoiding training stagnation.
$\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(x_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}))].$ (10)
The overall Privileged Loss $\mathcal{L}_{\mathcal{P}}$ is the expectation over these conditional losses:
$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right$.
$\quad\left.+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\right]$ (11)
where $h_{\text{DPO}}$ refers to the appropriate term from Eq. 8 or Eq. 9. The hyperparameter $\beta$ controls the preference strength.
SimPO-Inspired Loss: We also adapt insights from Simple Preference Optimization (SimPO) (Meng et al., 2024). Original SimPO for generative models uses length-normalized log-likelihoods and a target margin $\gamma$ in its preference scoring. Since our multi-label classification (MLC) setup deals with individual label scores $m(x_{i};\mathbf{w}_{t})$ where sequence length is inapplicable, we adapt SimPO’s core concept of a target margin $\gamma$ to our context. This results in a reference-free preference component that aims to ensure a minimum separation for preferred label scores, omitting length normalization but retaining the margin:
$L_{\text{pref}}^{\text{SimPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)-\gamma\right),$ (12)
where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$. The term $\beta\log(\cdot)$ captures the relative preference, while $\gamma$ enforces a desired separation margin. This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq. 11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq. 10) for $S_{il}=\emptyset$ remaining. More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
CPO-Inspired Loss: Adapting Contrastive Preference Optimization (Xu et al., 2024). If $S_{il}\neq\emptyset$, the preference component is:
$L_{\text{pref}}^{\text{CPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)\right).$ (13)
(using the same preferred/dispreferred logic as SimPO). CPO also includes an NLL regularizer. Thus, the overall CPO-inspired privileged loss, $\mathcal{L}_{\mathcal{P}}^{\text{CPO}}$, could be formulated by combining this preference term (when $S_{il}\neq\emptyset$) with a consistent NLL/BCE component for all true ground truth labels within $\mathcal{P}$:
[…]
where $k_{s}$ is a confusing example sampled from $S_{il}$, and $\lambda_{\text{CPO}}$ is a weighting hyperparameter. Note that the $\ell_{\text{BCE}}$ term is always applied, serving as both the NLL regularizer and the fallback for $S_{il}=\emptyset$ if the first term becomes zero. These alternatives offer different ways to model preferences. The non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq. 17) remains unchanged regardless of the privileged loss choice.
3.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
For non-privileged labels $j\in\bar{\mathcal{P}}$, our goal is to maintain performance relative to the reference model $\hat{\mathbf{w}}_{j}$, preventing degradation due to the focus on privileged labels. We use a standard base classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ (e.g., Binary Cross-Entropy) defined as:
$\ell(\mathbf{w}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\mathbf{w}_{j})+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}))],$ (15)
$\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\hat{\mathbf{w}}_{j})+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))].$ (16)
The Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$ employs a hinge mechanism, penalizing the model only if its loss $\ell(\mathbf{w}_{j})$ exceeds the reference model’s loss $\ell(\hat{\mathbf{w}}_{j})$ by more than a predefined slack margin $\epsilon\geq 0$:
$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
This ensures $\mathbf{w}_{j}$ is primarily updated only if performance on label $j$ drops significantly below the reference baseline plus $\epsilon$.
3.4 Group Robust Optimization Formulation
To effectively balance the objective for privileged labels ($\mathcal{L}_{\mathcal{P}}$) and the constraint for non-privileged labels ($\mathcal{L}_{\bar{\mathcal{P}}}$), we employ the Group Robust Preference Optimization (GRPO) framework (Ramesh et al., 2024). We define two distinct groups corresponding to our label partitions:
  • $G_{\mathcal{P}}$: Associated with the Privileged Loss $\mathcal{L}_{\mathcal{P}}$, focusing on triplets $(x_{i},l,k)$.
  • $G_{\bar{\mathcal{P}}}$: Associated with the Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$, focusing on pairs $(x_{i},j)$.
The FairPO objective becomes the GRPO minimax problem:
$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\,\geq\,0,\,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\,=\,1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})\right].$ (18)
Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss. The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$ that perform well even under this worst-case weighting. This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
3.5 Optimization Algorithm
The FairPO framework solves the minimax objective (Eq. 18) iteratively. A high-level overview is in Algorithm 1, with full details in Algorithm 2 (see Appendix B). In each training iteration, an instance and a label $r$ are sampled. The corresponding loss ($\mathcal{L}_{\mathcal{P}}^{(s)}$ or $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ via Eq. 11 or Eq. 17) and its gradient are computed based on whether $r$ is in the privileged set $\mathcal{P}$ or non-privileged set $\bar{\mathcal{P}}$.
The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step. A crucial aspect for the stability and effectiveness of this step is loss scaling. Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version. Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted $\bar{\mathcal{L}}_{g}^{\text{avg}}$. The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability. This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
This scaled loss, say $\Delta\tilde{\mathcal{L}}_{g}^{(s)}$, is then used in the exponential update: $\alpha_{g}^{\text{new}}\leftarrow\alpha_{g}^{(s)}\exp(\eta_{\alpha}\Delta\tilde{\mathcal{L}}_{g}^{(s)})$, followed by normalization so $\sum_{g}\alpha_{g}^{\text{new}}=1$. This scaling normalizes loss magnitudes across groups, makes weight updates sensitive to significant performance changes relative to recent history, and improves overall stability of the $\alpha$ dynamics. Subsequently, model parameters $\mathbf{w}^{(s+1)}$ are updated via mirror descent, using a gradient that combines group gradients weighted by the adaptively balanced $\alpha^{(s+1)}$.
Algorithm 1 FairPO Training Overview (DPO-inspired)
1:Initialize: Model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
3:For each training iteration $s=0,\dots,S-1$:
4:    Sample instance $(x_{i},y_{i})\sim\mathcal{D}$ and a label $r\in\mathcal{T}$.
5:    If $r\in\mathcal{P}$:
6:     Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (Eq. 11, conditionally using preference loss like Eq. 8/9 or BCE Eq. 10).
7:    Else if $r\in\bar{\mathcal{P}}$:
8:     Compute non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ for $(x_{i},r)$ (Eq. 17).
9:    Update group weights $\alpha^{(s+1)}$ via mirror ascent using $\mathcal{L}_{\mathcal{P}}^{(s)},\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ (GRPO step).
10:    Update model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
11:End For
12:Return Optimized parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(S)}$.
13:Full details are in Algorithm 2 (see Appendix B).
4 Experimental Setup
We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014 (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009). For our fairness setup, we define the privileged group ($\mathcal{P}$) as the 20% least frequent labels in each training set, with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$). Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021), where we fine-tune label-specific classifier heads on top of its features. The reference model parameters $\hat{\mathbf{w}}$ are obtained from a standard supervised fine-tuning (SFT) of this architecture with a Binary Cross-Entropy (BCE) loss.
Performance is assessed separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ sets using standard metrics: Mean Average Precision (mAP), Sample F1, Exact Match Ratio (EMR), and Accuracy. We compare FairPO against three baselines: (1) BCE-SFT, which also serves as our reference model; (2) BCE-SFT + Re-Weighting (RW), a simple loss up-weighting for $\mathcal{P}$ labels; and (3) Group DRO + BCE (Sagawa et al., 2020), which minimizes the worst-group loss. We test three variants of our framework, differing in their preference loss for privileged labels: FairPO-DPO (Eq. 11), FairPO-SimPO (Eq. 12), and FairPO-CPO (Eq. 14). Comprehensive details on datasets, preprocessing, model architecture, baseline implementations, and hyperparameter tuning are deferred to Appendices C, D, and E.
5 Results and Analysis
Table 1: Performance comparison on MS-COCO. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 86.32 90.65 61.43 64.89 94.89 98.12 35.78 36.98 Ref BCE SFT + RW 87.25 89.85 62.68 64.11 95.95 97.93 47.43 33.81 +0.93 GDRO + BCE 87.92 90.41 62.31 64.75 95.72 98.05 46.12 35.16 +1.60 FairPO-DPO 88.02 89.97 63.45 63.65 97.89 98.04 62.19 35.12 +1.70 FairPO-SimPO 87.67 88.76 62.34 63.12 95.69 97.78 45.32 32.34 +1.35 FairPO-CPO 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44
Table 2: Performance comparison on NUS-WIDE. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 63.53 70.24 48.12 55.83 91.51 95.22 19.32 11.56 Ref BCE SFT + RW 65.12 69.14 49.51 54.73 92.33 94.81 21.23 10.32 +1.59 GDRO + BCE 64.84 69.91 49.13 55.62 92.11 95.13 21.02 11.34 +1.31 FairPO-DPO 66.34 69.05 51.71 54.52 93.92 95.04 27.91 11.21 +2.81 FairPO-SimPO 64.11 68.03 48.82 53.81 91.94 94.52 20.18 10.19 +0.58 FairPO-CPO 67.12 69.83 52.21 55.24 94.31 95.12 31.87 11.25 +3.59
Tables 1 and 2 detail our findings. The standard BCE-SFT baseline confirms the fairness problem central to our motivation: on MS-COCO, a significant performance gap exists between the privileged ($\mathcal{P}$) labels (86.32% mAP) and the non-privileged ($\bar{\mathcal{P}}$) labels (90.65% mAP). This disparity highlights how standard aggregate losses can fail to ensure equitable performance for rare or difficult categories.
Our results demonstrate that FairPO variants, particularly FairPO-CPO, effectively address this challenge by navigating the fairness-performance trade-off in a controlled manner. On MS-COCO, FairPO-CPO achieves a substantial +3.44% mAP gain for $\mathcal{P}$ labels, while performance on $\bar{\mathcal{P}}$ labels dips by a negligible 0.31%. Similarly, on NUS-WIDE, it yields a +3.59% mAP gain for a minor 0.41% drop. This is not an uncontrolled side effect but a managed outcome, enforced by the GRPO mechanism and the constrained loss (Eq. 17), which together prevent significant harm to the non-privileged group while targeting improvements for the privileged one. By directly optimizing the model to rank true labels higher than their specific, dynamically identified confusing counterparts, FairPO sharpens discriminative power for the most challenging cases.
Among the variants, the reference-free FairPO-CPO proves most effective. This directly addresses a potential limitation of DPO-style methods, which can be sensitive to the quality of the reference model. FairPO-CPO’s strong performance, likely due to its dual objective of optimizing both relative preference and absolute correctness (via its NLL-like regularizer), demonstrates its robustness and versatility, making it a more practical choice than the reference-dependent FairPO-DPO or the fixed-margin FairPO-SimPO.
6 Ablation Studies
Table 3: Ablation on core components of FairPO-CPO (MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 55.20 37.15 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (-0.47) (-1.08) (+0.05) (-10.23) (+1.92)
Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT (86.32). Parentheses show change vs Full FairPO-CPO.
Preference Detail Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 (Conf. Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 22.54 35.10 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) (-42.89) (-0.13) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 60.75 34.90 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07) (-4.68) (-0.33)
We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions (Tables 3 and 4). Our analysis of core components (Table 3) confirms the criticality of each part. Replacing the preference loss with BCE reduces privileged gains, removing the $\bar{\mathcal{P}}$ constraint harms non-privileged performance, and using fixed weights instead of GRPO degrades both groups. To further isolate the value of our fairness framework, we tested a Global CPO variant that applies the preference objective to all labels without the privileged/non-privileged split or GRPO. While this proves the preference objective is a strong general-purpose loss (achieving a +2.23% mAP gain over BCE), the full FairPO framework is decisively superior for the targeted fairness task. By applying CPO globally, the performance gains are unfocused, resulting in a 1.21 point lower mAP on the privileged set. This demonstrates that the full FairPO architecture is crucial for focusing the model’s capacity to resolve the hardest discrimination challenges within the privileged group.
Furthermore, the design of the preference objective itself is crucial (Table 4). Restricting it to only Confusing Negatives (for $y_{il}=+1$) causes a profound performance drop, as this neglects the far more common scenario of ranking true negatives below confusing positives for rare labels, making the learning signal exceptionally sparse. Similarly, removing the BCE Fallback for non-confusing instances degrades performance, proving that a standard classification signal on easier cases is vital for model stability. Both aspects of our full formulation are thus essential, demonstrating that while targeted preference optimization is powerful, it must be complemented by standard losses and robust balancing for stable, effective training.
7 Related Work
Recent efforts in fair MLC address complex challenges like label imbalance impacting tail labels (Guo and Wang, 2021), subjective fairness (Liu et al., 2023), and class-incremental learning (Dong et al., 2025). FairPO contributes a novel approach by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets. It applies distinct, fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and manages the trade-off with a robustness framework, differentiating our targeted approach from prior work.
We adapt recent advances in preference optimization, originally developed for aligning LLMs (Ouyang and Others, 2022; Christiano et al., 2023). Techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and its reference-free variants CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies from preference pairs. Rather than ranking entire outputs, we repurpose these methods to specifically differentiate true label scores from their dynamically identified confusing counterparts within the privileged set, thereby sharpening critical decision boundaries. To balance our objectives, we employ a Group Robust Optimization strategy inspired by Group DRO (Sagawa et al., 2020; Rice et al., 2021) and GRPO (Ramesh et al., 2024). While these methods typically balance performance across data or preference groups, FairPO uniquely defines its groups by our label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$). It then uses GRPO’s adaptive weighting to balance their distinct, custom-formulated loss objectives, providing a principled mechanism for managing the specific fairness-performance trade-offs in our MLC context.
8 Discussion
In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification. Our experiments, particularly with the FairPO-CPO variant, highlight the value of nuanced preference signals for navigating complex fairness-performance trade-offs in challenging discrimination tasks. While promising, FairPO has technical limitations: its dynamic confusing set can lead to instability or sparse signals for rare labels; the DPO-based variant and non-privileged constraint rely on a well-calibrated reference model; and GRPO’s balancing of heterogeneous losses requires careful tuning and lacks theoretical convergence guarantees. These challenges also define our future work, which includes extending FairPO’s principles to multi-label attribute generation (Appendix A), conducting comprehensive empirical validation on more datasets and modalities, analyzing alternative label partitioning strategies, and pursuing theoretical insights into the framework’s convergence properties.
Appendix A Adapting FairPO for Multi-Attribute Generation
This section outlines our planned extension of FairPO to multi-attribute generation, a conceptual direction for future work. The goal is to generate a sequence $y$ from a prompt $x$ using a policy $\pi_{\mathbf{w}}(y|x)$ that aligns with fairness goals over a set of attributes $\mathcal{A}$. The core idea involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets and retaining the GRPO minimax structure (Eq. 18). The group losses would be defined over a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$, with preference losses like DPO applied to the log-probabilities of entire generated sequences rather than individual label scores.
Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For privileged attributes $j\in\mathcal{P}$, the goal is to ensure the learned policy $\pi_{\mathbf{w}}$ strongly reflects preferences $y_{w}\succ y_{l}$ established by that attribute. This is achieved using a standard DPO loss, averaged over the privileged subset of the preference data:
$\mathcal{L}_{\mathcal{P}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},j)\sim\mathcal{D}_{pref}\mid j\in\mathcal{P}}\left[-\log\sigma\left(\beta\cdot h_{\pi_{\mathbf{w}}}(x,y_{w},y_{l})\right)\right]$ (19)
Minimizing this loss directly encourages the model to favor preferred sequences for preferences driven by privileged attributes, relative to the reference policy $\pi_{\text{ref}}$.
Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): For non-privileged attributes $k\in\bar{\mathcal{P}}$, the objective remains analogous to the classification setting: preventing significant performance degradation. This is accomplished with a hinge formulation based on the DPO loss:
$\mathcal{L}_{\bar{\mathcal{P}}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},k)\sim\mathcal{D}_{pref}\mid k\in\bar{\mathcal{P}}}\left[\max\left(0,\mathcal{L}_{DPO}(\pi_{\mathbf{w}},\pi_{\text{ref}};x,y_{w},y_{l})-(\log 2)-\epsilon^{\prime}\right)\right].$ (20)
This penalizes the model only if its preference modeling for non-privileged attributes degrades substantially beyond baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$. The overall FairPO objective would then use GRPO to balance these two losses.
Planned Experiments: The planned task is generating text from a prompt, requiring preference datasets where the driving attribute $j_{i}$ is known. This could involve partitioning attributes like ‘Helpfulness’ or ‘Harmlessness’ from existing RLHF datasets (e.g., Anthropic HH-RLHF [Bai and Others, 2022], Summarization preferences [Stiennon et al., 2022]) or creating new, explicitly annotated data. We plan to fine-tune pretrained language models (e.g., Gemma [Team, 2024], Llama [Touvron and Others, 2023], Mistral [Jiang and Others, 2023]) with this objective. Evaluation would rely on preference-based metrics, such as win rates for privileged and non-privileged attribute groups, compared against baselines like standard DPO and SFT. Future experiments would also explore hyperparameter effects and adapting the SimPO/CPO loss formulations for generation.
Appendix B FairPO Algorithm
The FairPO framework is trained iteratively to solve the minimax objective presented in Eq. 18. The detailed procedure, which is inspired by the DPO-based variant of FairPO, is provided in Algorithm  2.
Initialization: The training process begins by initializing the model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$, for instance by copying them from a pre-trained reference model $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$. The adaptive group weights, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$, are typically set to uniform values such as 0.5 each.
Iterative Training Loop: The core of the framework is an iterative training loop. In each step, an instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$, and a single label $r\in\mathcal{T}$ is randomly selected from that instance for processing. The subsequent steps depend on whether this sampled label belongs to the privileged or non-privileged set.
If the sampled label $r$ is in the privileged set $\mathcal{P}$, the algorithm first identifies if a confusing set $S_{il}$ exists for that label (where $l=r$), as detailed in Algorithm 2. The loss computation is then conditional on this set:
  • If confusing examples exist ($S_{il}\neq\emptyset$), a DPO-inspired preference loss is computed between label $l$ and a randomly sampled confusing example $k\in S_{il}$. This preference loss directly encourages the model to improve its ranking of $l$ relative to its specific confounder $k$.
  • If no confusing examples are found ($S_{il}=\emptyset$), the algorithm reverts to a standard base classification loss (e.g., BCE, Eq. 10) for label $l$. This fallback is crucial as it ensures the model continues to receive a learning signal on easier instances, promoting stable training.
The loss calculated from either of these cases contributes to the current step’s privileged group loss, $\mathcal{L}_{\mathcal{P}}^{(s)}$. Conversely, if the sampled label $r$ belongs to the non-privileged set $\bar{\mathcal{P}}$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed according to Eq. 17. This loss penalizes the model only if its performance on the label $j=r$ deviates from the reference model’s performance by more than a predefined slack margin $\epsilon$.
After the appropriate group loss is computed, the GRPO mechanism performs two key updates. First, the Adaptive Weight Update adjusts the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ using a mirror ascent step. This step uses an exponential weighting based on the current (and scaled) group losses, dynamically increasing the focus on the group that is currently performing worse (Lines 39-41). Second, the Model Parameter Update updates all model parameters $\mathbf{w}_{t}$ via a mirror descent step, using a combined gradient that is weighted by the newly updated adaptive weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$.
This entire process repeats for a predefined number of iterations or until convergence, allowing FairPO to dynamically balance its objectives to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq. 12 or 13). The overall GRPO structure and non-privileged handling are consistent across all variants.
Algorithm 2 FairPO Algorithm for Multi-Label Classification (DPO-inspired)
1:Initialize: $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
2:Choose: $\eta_{\mathbf{w}},\eta_{\alpha}$, $\beta$, $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, $\epsilon$.
3:for $s=0$ to $S$ (MaxIterations) do
4:    Sample an example: $(x_{i},[y_{i1},\dots,y_{iT}])\in\mathcal{D}\sim p_{\mathcal{D}}(.)$.
5:    Initialize group losses for this step: $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow 0$, $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow 0$.
6:    Initialize gradients: $g^{t}_{\mathcal{P}}\leftarrow\vec{0}$, $g^{t}_{\bar{\mathcal{P}}}\leftarrow\vec{0}\quad\forall t\in\mathcal{T}$.
7:    Forward pass: $m(x_{i};\mathbf{w}_{t}^{(s)})\leftarrow\sigma(\mathbf{w}_{t}^{(s)^{T}}\mathbf{z}_{i})$ where $\mathbf{z}_{i}\leftarrow\pi_{\theta}(x_{i})\quad\forall t\in\mathcal{T}$.
8:    Sample a label: $r\in\mathcal{T}\sim Uniform(\frac{1}{|\mathcal{T}|})$.
9:    if $r\in\mathcal{P}$ then $\triangleright$ Handle privileged label
10:     $l\leftarrow r$, $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
11:     if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
12:      $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{neg}}$
13:     else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
14:      $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{pos}}$
15:     end if
16:     if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
17:      Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
18:      if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
19:        $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
20:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
21:      else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
22:        $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
23:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
24:      end if
25:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
26:     else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
27:      $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
28:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
29:     end if
30:    else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
31:     $j\leftarrow r$
32:     $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
33:     $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
34:     $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$, $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
35:    end if
36:    $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$ $\triangleright$ Mirror ascent
37:    $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$, $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$ $\triangleright$ Weight normalization
38:    $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}\triangleright$ Mirror descent
39:end for
40:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$
Appendix C Dataset and Preprocessing Details
MS-COCO 2014 [Lin et al., 2015]: We used the official 2014 train/val splits. The training set contains 82,783 images and the validation set (used as our test set) contains 40,504 images. There are 80 object categories. The privileged set $\mathcal{P}$ consisted of the 16 labels (20% of 80) with the lowest frequency in the training set. The remaining 64 labels formed $\bar{\mathcal{P}}$.
NUS-WIDE [Chua et al., 2009]: This dataset contains 269,648 images with 81 concept labels. We used the common split of 161,789 images for training and 107,859 for testing. The privileged set $\mathcal{P}$ consisted of the 16 labels (approx. 20% of 81) with the lowest frequency in the training set. The remaining 65 labels formed $\bar{\mathcal{P}}$.
Image Preprocessing: For both datasets, images were resized to $224\times 224$ pixels and normalized using the standard ImageNet mean and standard deviation, consistent with the ViT pretraining. Standard data augmentations like random horizontal flips and random resized crops were applied during training.
Appendix D Baseline Experimental Details
This section provides further details on the implementation and hyperparameter tuning for the baseline methods used in our experiments. Unless otherwise specified, all baselines were trained using the same base Vision Transformer (ViT) architecture, optimizer (AdamW), number of epochs, and batch size as the main FairPO experiments for fair comparison. Hyperparameters specific to each baseline were tuned on the validation set of MS-COCO and NUS-WIDE. The privileged group $\mathcal{P}$ comprised the 20% least frequent labels, and the non-privileged group $\bar{\mathcal{P}}$ comprised the remaining 80%.
D.1 BCE-SFT (Reference Model)
The BCE-SFT baseline represents standard supervised fine-tuning. Its objective is to minimize the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels, where the loss for an instance $(x_{i},y_{i})$ is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$. For implementation, a separate non-linear MLP classifier head is trained for each label $t$ on top of frozen ViT features, consistent with the main FairPO architecture. The primary hyperparameter tuned was the learning rate for these classifier heads, selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ to achieve the best overall mean Average Precision (mAP) on the validation set. Weight decay was set to 0.01, matching the FairPO experiments.
This model serves two critical roles in our study. First, it acts as a direct performance baseline against which we measure fairness improvements. Second, the final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used throughout the FairPO framework, both for the DPO-based losses and for the non-privileged loss constraint. The model was trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
D.2 BCE-SFT + Privileged Re-Weighting (RW)
This baseline modifies the standard BCE-SFT by applying static loss re-weighting to improve performance on the privileged labels $\mathcal{P}$. The objective is to assign a higher weight to their contribution in the total loss function:
$\mathcal{L}_{\text{RW}}=\lambda_{\mathcal{P}}\sum_{l\in\mathcal{P}}\text{BCE}(m(x_{i};\mathbf{w}_{l}),y_{il})+\lambda_{\bar{\mathcal{P}}}\sum_{j\in\bar{\mathcal{P}}}\text{BCE}(m(x_{i};\mathbf{w}_{j}),y_{ij})$
For this setup, we use the same per-label classifier architecture as BCE-SFT, but the loss computation is modified to incorporate the weights. The key hyperparameter is the privileged group weight $\lambda_{\mathcal{P}}$, which was tuned via grid search from values $\{2,3,5,8,10\}$. The value was chosen to maximize mAP on the privileged group without an excessive drop in performance on the non-privileged group. For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective, while $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$. Training was conducted similarly to BCE-SFT.
D.3 Group DRO + BCE
This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss. The objective is to minimize the worst-case BCE loss across the predefined privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) groups:
$\min_{\{\mathbf{w}_{t}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\geq 0,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\text{BCE}}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})\right]$
Here, $\mathcal{L}_{\text{BCE}}(\mathcal{G})$ is the average BCE loss for all labels within a group $\mathcal{G}$. The implementation uses the same BCE-SFT architecture. Per-label BCE losses are computed and then aggregated separately for the $\mathcal{P}$ and $\bar{\mathcal{P}}$ groups. Model parameters $\{\mathbf{w}_{t}\}$ are updated with AdamW, while the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step: $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$. The final loss for backpropagation is the adaptively weighted sum of the group losses.
For hyperparameters, the model learning rate was tuned as for BCE-SFT. The group weight learning rate $\eta_{\alpha}$ was tuned from $\{0.01,0.05,0.1,0.2\}$, with $\eta_{\alpha}=0.05$ selected for its stable convergence and good worst-group performance. For the reported results, raw group losses were used in the exponential update, though we note that a loss scaling approach similar to that in FairPO could also be employed for enhanced stability. The training protocol regarding epochs and batching was identical to FairPO.
Appendix E FairPO Experimental Details
E.1 Common Setup for All FairPO Variants
Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison. The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k. During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features. All experiments were conducted on the MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009] datasets. Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops. The AdamW optimizer [Loshchilov and Hutter, 2019] was used to update all trainable parameters. Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set. The reference model parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, required for FairPO-DPO and the non-privileged loss constraint in all variants, were obtained from a BCE-SFT model detailed in Appendix D.
E.2 Per-Label Non-Linear MLP Classifier Head
For each of the $T$ labels in a dataset, we employed a dedicated and independent non-linear Multi-Layer Perceptron (MLP) head to predict the probability of that label being positive. Using separate MLP heads allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics, which is particularly beneficial for labels with varying difficulty. Each MLP head takes the $d$-dimensional feature vector (where $d=768$ for ViT-Base) from the ViT’s [CLS] token as input and outputs a single logit. The final probability score $m(x_{i};\mathbf{w}_{t})$ is obtained by applying a sigmoid function to this logit. The specific architecture for each MLP head is as follows:
  1. 1.
    Linear Layer: $d\rightarrow 256$ neurons, followed by ReLU Activation
  2. 2.
    Linear Layer: $256\rightarrow 64$ neurons, followed by ReLU Activation
  3. 3.
    Linear Layer: $64\rightarrow 16$ neurons, followed by ReLU Activation
  4. 4.
    Linear Layer: $16\rightarrow 4$ neurons, followed by ReLU Activation
  5. 5.
    Linear Layer (Output): $4\rightarrow 1$ neuron (producing the logit)
The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are unique to that label. All parameters within these MLP heads were fully trainable during both the SFT pre-training (for the reference model) and the final FairPO fine-tuning.
E.3 FairPO Variant-Specific Hyperparameters
The core FairPO learning rates, $\eta_{\mathbf{w}}$ for model parameters and $\eta_{\alpha}$ for GRPO weights, were tuned via grid search. For $\eta_{\mathbf{w}}$ (the AdamW LR), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$, and for $\eta_{\alpha}$, values were explored from $\{0.01,0.05,0.1,0.2\}$. After tuning, the final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$. The specific hyperparameters for each FairPO variant’s privileged loss component (e.g., $\beta$, $\gamma$, $\lambda_{\text{CPO}}$) were also tuned on the designated validation set for each dataset, with final values reported in Table 5 (Appendix F). Throughout all experiments, the GRPO mechanism with scaled loss updates, as described in Section 3, was used to adaptively balance the privileged and non-privileged loss terms.
Appendix F Hyperparameter Details
F.1 Hyperparameters Tuning
For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019] with a batch size of 32. The base model was a vit-base-patch16-224222https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k. The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$, while the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$. All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP, using a patience of 5 epochs [^3] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
Specific hyperparameters for each FairPO variant were tuned via grid search on the validation set. For FairPO-DPO, the strength parameter $\beta$ was chosen from $\{0.1,0.5,1.0\}$, and the non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$. For FairPO-SimPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the margin $\gamma$ from $\{0.05,0.1,0.2\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. For FairPO-CPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the NLL regularizer weight $\lambda_{\text{CPO}}$ from $\{0.1,0.5,1.0\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
For our baselines, the loss weight for privileged labels in BCE-SFT + Privileged Re-Weighting was set to 5 after tuning from $\{2,3,5,10\}$. The group weight learning rate $\eta_{\alpha}$ for GDRO + BCE was tuned similarly to FairPO.
F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE. Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$. Base learning rate $\eta_{\mathbf{w}}$ was 1e41e-4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
Dataset Method $\beta$ / $\lambda_{\text{CPO}}$ $\gamma$ (SimPO only) $\epsilon$
MS-COCO FairPO-DPO $\beta=0.5$ N/A $0.05$
FairPO-SimPO $\beta=0.5$ $0.1$ $0.05$
FairPO-CPO $\beta=0.5,\lambda_{\text{CPO}}=0.5$ N/A $0.01$
NUS-WIDE FairPO-DPO $\beta=0.1$ N/A $0.05$
FairPO-SimPO $\beta=0.1$ $0.05$ $0.1$
FairPO-CPO $\beta=0.1,\lambda_{\text{CPO}}=0.5$ N/A $0.05$
Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values. Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results.
Table 6: Sensitivity of FairPO-CPO mAP on MS-COCO to preference strength $\beta$ (with $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$).
$\beta$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.05 88.95 90.28
0.1 89.32 90.31
0.5 (Optimal) 89.76 90.34
1.0 89.51 90.25
2.0 88.67 90.11
Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed. A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
Table 7: Sensitivity of FairPO-CPO mAP on MS-COCO to non-privileged slack $\epsilon$ (with $\beta=0.5,\lambda_{\text{CPO}}=0.5$).
$\epsilon$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.001 89.65 90.40
0.01 (Optimal) 89.76 90.34
0.05 89.81 90.22
0.1 89.70 90.05
0.2 89.55 89.87